### v6

@@ 3.6 Discussion of the Main Assumptions
Players can agree to disagree over the true model of the game. In this case, Ann sees Bob’s model as misspecified and uninformative about $Q$. The test LR implies that players ignore evidence against their model, unless surprised. [^22] The likelihood ratio test was popularized by (Neyman et al.,, 1933). Variations of this rule have been adopted in many forms in the economic literature (e.g., Hong et al.,, 2007; Ortoleva,, 2012; Ba,, 2022). This captures two largely documented behavioral biases: overconfidence in own worldviews (Evans,, 1990; Nickerson,, 1998; Andreoni and Mylovanov,, 2012), and over-reaction to low probability events (for a review, Hong et al.,, 2007; Ortoleva,, 2012). Finally, players only consider alternative models held by at least one group member. If they disagree, the assumption captures disagreement-induced binary thinking (Lewis et al.,, 2019). If they are like-minded, it captures groupthink, inhibiting their ability to change perspectives (Janis,, 1982). [^23] Our qualitative results could be replicated in a Bayesian framework under additional mild assumptions. However, our driving force would be harder to isolate due to players’ own learning incentives. Most importantly, we believe that the model change approach of this paper better matches our motivating examples, while also qualitatively capturing behavioral biases shown to arise when different worldviews collide.
@@ 4 One Technology: The Power of Skepticism
The lemma is proven in the appendix. It has a simple interpretation: both Ann and Bob believe that by working harder they will prove to their disagreeing counterpart that they hold the true model $m^{i}$. This is always true subjectively, regardless of the true model $Q$ [^24] The result follows from our assumption on effort informativeness and known results by Neyman et al., (1933) and Blackwell and Girshick, (1962) The result is at the core of the persuasion incentives highlighted in the paper: players believe that more productive actions – higher effort – will be more likely to trigger a model change.
@@
The intuition for this result is as follows. First, Ann and Bob anticipate that if they exert more effort, they will produce more information about the production technology. If they disagree, they expect this to falsify the model of the other group member more often, by lemma 4.1. Next, note that if Bob switches to model $\mathcal{H}$, he will work harder in the second period, because $e^{\mathcal{H}}>e^{\mathcal{L}}$. This benefits Ann through the production externality. Conversely, when Ann is the one changing her mind after the first period – switching to $\mathcal{L}$ – her effort and expected output in the second period will decrease, lowering Bob’s future expected payoff. As a consequence, Ann works more in the first period than she would if Bob was like-minded, in order to “bring him on board.” Bob, in contrast, might be induced to reduce effort to prevent Ann from becoming skeptical. However, when $e^{\mathcal{L}}=0$, as in section 2, Bob’s effort does not decrease under disagreement, as he’s not engaging in production in the first place.
@@
Assume that $\mathcal{L}$ discourages effort, $e^{\mathcal{L}}=0$. The following holds:
  1. (i)
    If the optimist holds the true model, a disagreeing team is expected to produce more than the average of the two like-minded teams. If the skeptic holds the true model, a disagreeing team can be expected to produce more or less than the average of the two like-minded teams.
@@
For the next proposition, assume that $b$ is large enough that there exists $\hat{e}_{Q}\in\mathcal{E}$ such that $\mathbb{E}_{Q}[Y|\hat{e}_{Q}]=4\mathbb{E}_{Q}[Y|e^{\mathcal{H}}]$. Let $\Delta=\mathbb{E}_{\mathcal{H}}[v(Y)|e^{\mathcal{H}}]-\mathbb{E}_{\mathcal{H}}[v(Y)|e^{\mathcal{L}}]$.
@@
Proposition 4.2 suggests that adding to a team of optimists a skeptic (who doesn’t exert effort unless convinced) could – under some circumstances – be more valuable to an output-maximizing team manager than adding an additional optimist: the presence of a skeptic allows the manager to get the most out of the more optimist team members, those already convinced that the effort will pay off. For the benefit of disagreement to materialize, the externalities must be sufficiently high. If Ann expects to strongly benefit from changing Bob’s mind, she will increase her effort and first-period output considerably, possibly producing more than a group composed of individually most productive types. If such externalities are low, Ann will not care about Bob’s view and behave similarly to a myopic player.
@@ 5 Competition of Ideas
We now analyze the problem with $|K|=2$, $K=\{x,y\}$. [^25] The extension of our main results to $|K|>2$ is trivial given our focus on two-member group and the assumption that each group member only switch to a model adopted by someone in a previous stage of the game, as discussed in section 3.6. With two technologies, the set of models becomes $M=\{(\mathcal{H}_{x},\mathcal{H}_{y}),(\mathcal{H}_{x},\mathcal{L}_{y}),(\mathcal{L}_{x},\mathcal{H}_{y}),(\mathcal{L}_{x},\mathcal{L}_{y})\}$, where subscripts are only used to keep track of the views corresponding to each technology.
The multi-technology setup admits many possible interpretations. For instance, our group could consist of board members of a company trying to decide which business project to prioritize between a number of promising, but uncertain alternatives. They could be investors deciding between two competing investment portfolios. They could be academics disagreeing over which research method, or theory, will yield the best answers to a research question. They might be collaborating artists, disagreeing over the best way to create an artwork or song. In all cases, Ann and Bob have access to alternative methods of production. Importantly, we maintain the assumption that the players’ interests are aligned, to some degree: each of them benefits if the other adopts the best production method – no matter how strong the disagreement over what such method is. [^26] For instance, if Ann and Bob are two scientists facing a similar research question, we assume that they share the common goal of pushing knowledge as far as possible in their field.
With two technologies, disagreement can be of two kinds. As in section 4, it can be “vertical,” when players differ in their degree of optimism. [^27] This is the case, for instance, if Ann holds $m^{A}=(\mathcal{H}_{x},\mathcal{H}_{y})$ while Bob holds $m^{B}=(\mathcal{L}_{x},\mathcal{L}_{y})$, or if $m^{A}=(\mathcal{H}_{x},\mathcal{H}_{y})$ and $m^{B}=(\mathcal{H}_{x},\mathcal{L}_{y})$. But it can also be “horizontal,” when they are optimistic about different technologies – so that they cannot be ranked in terms of their optimism. For instance, this happens if $m^{A}=(\mathcal{H}_{x},\mathcal{L}_{y})$ and $m^{B}=(\mathcal{L}_{x},\mathcal{H}_{y})$.
@@
The intuition for the result is as follows. If Ann and Bob start the game like-minded, in the best case scenario they will exert effort $e^{\mathcal{H}}$ in every period – this happens provided that they are optimistic about at least one technology. With initial horizontal disagreement, for any technology assignment, players are motivated to exert more effort in $t=1$ than statically optimal given the corresponding technology view. By working harder, they increase the probability of their colleague’s model change, which, they believe, would ensure that the latter adopts the best technology in $t=2$. Hence, when initially assigned to the technology they are optimistic about, both Ann and Bob choose effort levels above $e^{\mathcal{H}}$. When ($Q_{x}=Q_{y}$), the increase in effort always translates into higher expected output.
@@
    If the opposing models $(\mathcal{H}_{x},\mathcal{L}_{y})$ and $(\mathcal{L}_{x},\mathcal{H}_{y})$ are equally likely to be correct, a team with horizontal disagreement performs on average better than any like-minded team. Formally, if $Q=(\mathcal{H}_{x},\mathcal{L}_{y})$ with probability $p(\mathcal{H}_{x},\mathcal{L}_{y})=\frac{1}{2}$ and $Q=(\mathcal{L}_{x},\mathcal{H}_{y})$ with probability $p(\mathcal{L}_{x},\mathcal{H}_{y})=\frac{1}{2}$, then
    $\mathbb{E}_{p}\left[\hat{Y}((\mathcal{H}_{x},\mathcal{L}_{y}),(\mathcal{L}_{x},\mathcal{H}_{y}),Q)\right]\geq\mathbb{E}_{p}\left[\hat{Y}(m,m,Q)\right]\quad\forall m\in M.$
@@
A manager wants to form a team of two engineers to develop a new product. The product can be based on two very different innovative technologies. Some of the engineers of the firm are optimistic about both alternatives (i.e., hold view $\mathcal{H}$ for both technologies), others support only one. The manager knows that typically one approach proves better than the other – one technology is of type $\mathcal{H}$ and the other of type $\mathcal{L}$. – but, from her point of view, both technologies are equally promising ex-ante. Proposition 5.2 (ii) tells us that the manager will maximize the expected output of the team if she picks co-workers with opposite views, $(\mathcal{H},\mathcal{L})$ vs $(\mathcal{L},\mathcal{H})$, provided that $\mathcal{H}$ and $\mathcal{L}$ are equally falsifiable. The benefit of disagreement is twofold: First, both team members will try harder to make more progress during the first production period. Such progress is valuable on its own. Second, the higher first-period effort conveys additional information about the promise of the two technologies, making it more likely that the team will adopt the best technology in the second period.
The takeaway of this section is simple and yet, we believe, important and non-trivial. When two similarly good production technologies are available, disagreement on the best way to produce might increase group production. We focused on a mechanism that builds on the idea of different perspectives and shows that these differences can be useful even when they do not complement each other, that is, when there is competition of ideas. When they have the common goal of increasing each others’ output, disagreeing people will challenge each other, and work harder to prove, with their successes, that their perspective is valid, beneficial, and worth adopting. The resulting increase in group output is where – we believe – lies the “disagreement dividend.”
@@ 6 Informativeness, Externalities, and Welfare
A similarly crucial role is played by the assumption that each team member is ceteris paribus better off when other members produce more. The positive externality creates the incentive to persuade disagreeing co-workers to abandon skeptical views, as well as the one to convince them to adopt the best technology. Clearly, if other players’ output entered utility negatively, the incentive would be very different. To see this, consider equations 1, 2 and 3 of our illustrative example of section 2, but let $\beta<0$ – so that externalities are negative. By inspecting the equations, it is easy to see that $\beta<0$ implies that the value of triggering a co-worker’s change in mind at the end of the first period becomes negative: as a result, disagreement reduces Ann’s first-period effort and expected output, rather than increasing them.
@@
We omit the formal proofs for the claim, as such proofs would follow very closely the ones reported in the appendix. Instead, we provide the intuition for the one-technology case (the multiple-technology case follows the same intuition). Let the alternative assumptions of claim 1 hold, and consider a disagreeing team. Note that the player with view $\mathcal{H}$ will be hurt from a change of mind of the $\mathcal{L}$-view player: by assumption 2, the skeptic works harder after becoming optimist, damaging her co-worker through the negative production externalities of point (i). This force pushes the optimistic player to reduce information arrival in the first period, in order to decrease the (subjective) probability of a co-worker’s change in mind. However, when the relation between effort and informativeness is inverse – as of point (ii) – the way to reduce information arrival is by working harder. Hence, even under these alternative assumptions, disagreement pushes optimists to work harder and, if negative externalities are strong enough, a disagreeing team will produce on average more than any like-minded team. [^29] A few details of the propositions will indeed need to be intuitively modified for the result to hold under the assumptions of claim 1. In particular, the second part of proposition 4.2 (ii) holds for $\Delta$ negative enough; and the requirement for inequalities to be strict becomes that $\phi^{m^{-i}}_{m^{i}}(\mathbf{e},\mathbf{k}|m^{i})$ is strictly decreasing in effort. Regardless of this equivalence, we believe that our original specification – with positive externalities and more information arriving the more a team works on a project – is particularly realistic and captures well the zest of most of our motivating examples.
We conclude the discussion with a word on the welfare implications of disagreement. A welfare analysis would be particularly complex in our setup, which does not impose strong assumptions on the true process $Q$. Without such assumptions, it is hard to tell, for instance, whether the effort levels that maximize team members’ joint expected payoffs (or are Pareto efficient in terms of expected utility) are above or below $e^{\mathcal{H}}$ and $e^{\mathcal{L}}$. What we can say with certainty is that if the assumptions of proposition 5.2 part (i) hold and if, additionally, $Q_{x}=Q_{y}=\mathcal{H}$, any Pareto efficient stage effort must be above $e^{\mathcal{H}}$. Hence the boost in players’ effort generated by horizontal disagreement can – as in the example presented in section 2 – lead to a Pareto improvement. Not much can be concluded, however, in general. [^30] For instance, note that with $|K|=1$ and $Q=\mathcal{H}$, the expected payoff of an optimist is always higher when she’s paired with another optimist. In such case, disagreement cannot lead to a Pareto improvement in terms of expected payoffs.

### v7

@@ 3.6 Discussion of the Main Assumptions
Players can agree to disagree over the true model. In this case, Ann sees Bob’s model as misspecified and uninformative about $Q$. The test LR implies that players ignore evidence against their model, unless surprised. [^23] The likelihood ratio test was popularized by (Neyman et al.,, 1933). Variations of this rule have been adopted in many forms in the economic literature (e.g., Hong et al.,, 2007; Ortoleva,, 2012; Ba,, 2022). This captures two largely documented behavioral biases: overconfidence in own worldviews (e.g., Evans,, 1990; Nickerson,, 1998; Andreoni and Mylovanov,, 2012; Galperti,, 2019, for a review), and over-reaction to low probability events (for a review, Hong et al.,, 2007; Ortoleva,, 2012). Finally, players only consider alternative models held by other group members. If they disagree, this captures disagreement-induced binary thinking (Lewis et al.,, 2019). If they are like-minded, it captures groupthink, inhibiting their ability to change perspectives (Janis,, 1982). [^24] Our qualitative results could be replicated in a Bayesian framework under additional assumptions. However, our driving force would be harder to isolate due to players’ own learning incentives. Most importantly, we believe that the model change approach of this paper better matches our motivating examples, while also qualitatively capturing behavioral biases shown to arise when different worldviews collide.
@@ 4 One Technology: The Power of Skepticism
The lemma is proven in the appendix. It has a simple interpretation: both Ann and Bob believe that by working harder they will prove to their disagreeing counterpart that they hold the true model $m^{i}$. This is always true subjectively, regardless of the true model $Q$ [^25] The result follows from our assumption on effort informativeness and known results by Neyman et al., (1933) and Blackwell and Girshick, (1962) The result is at the core of the persuasion incentives highlighted in the paper: players believe that more productive actions — higher effort — will be more likely to trigger a model change. This force is akin to the “information validates the prior” finding of Kartik et al., (2021) in a Bayesian setting. [^26] Kartik et al., (2021) shows that, under some ordering conditions, if Ann and Bob have different priors, Ann expects more information to move Bob’s posterior mean closer to her prior mean.
@@
The intuition for this result is as follows. First, Ann and Bob anticipate that if they exert more effort, they will produce more information about the production technology. If they disagree, they expect this to falsify the model of the other group member more often, by lemma 4.1. Next, note that if Bob switches to model $\mathcal{H}$, he will work harder in the second period, because $e^{\mathcal{H}}>e^{\mathcal{L}}$. This benefits Ann through the production externality. Conversely, when Ann is the one changing her mind after the first period — switching to $\mathcal{L}$ — her effort and expected output in the second period will decrease, lowering Bob’s future expected payoff. As a consequence, Ann works more in the first period than she would if Bob was like-minded, in order to “bring him on board.” Bob, in contrast, might be induced to reduce effort to prevent Ann from becoming skeptical. However, when $e^{\mathcal{L}}=0$, as in section 2, Bob’s effort does not decrease under disagreement, as he’s not engaging in production in the first place.
@@
Assume that $\mathcal{L}$ discourages effort, $e^{\mathcal{L}}=0$. Then, if the optimist holds the true model, a disagreeing team is expected to produce more than the average of the two like-minded teams. If the skeptic holds the true model, a disagreeing team can be expected to produce more or less than the average of the two like-minded teams.
@@
For the next proposition, assume that $b$ (the upper bound on effort) is large enough that there exists $\hat{e}_{Q}\in\mathcal{E}$ such that $\mathbb{E}_{Q}[Y|\hat{e}_{Q}]=4\mathbb{E}_{Q}[Y|e^{\mathcal{H}}]$. Let $\Delta=\mathbb{E}_{\mathcal{H}}[v(Y)|e^{\mathcal{H}}]-\mathbb{E}_{\mathcal{H}}[v(Y)|e^{\mathcal{L}}]$.
@@
Proposition 4.2 suggests that adding to a team of optimists a skeptic (who doesn’t exert effort unless convinced) could — under some circumstances — be more valuable to an output-maximizing team manager than adding an additional optimist: the presence of a skeptic allows the manager to get the most out of the more optimist team members, those already convinced that the effort will pay off. For the benefit of disagreement to materialize, the externalities must be sufficiently high. If Ann expects to strongly benefit from changing Bob’s mind, she will increase her effort and first-period output considerably, possibly producing more than a group composed of individually most productive types. If such externalities are low, Ann will not care about Bob’s view and behave similarly to a myopic player.
@@ 5 Competition of Ideas
We now analyze the problem with $|K|=2$, $K=\{x,y\}$. [^27] The extension of our main results to $|K|>2$ is trivial given our focus on two-member groups and the assumption that each group member only switches to a model adopted by someone in a previous stage of the game, as discussed in section 3.6. Consistently with the conditions of propositions 5.1 and 5.2, an immediate extension requires that there exist $x,y\in K,x\neq y,$ such that (i) $Q_{x}=Q_{y}$, and (ii) $Q_{x}(\cdot|e)$ weakly first-order stochastically dominates $Q_{k}(\cdot|e)$ for any $e\in\mathcal{E}$ and $k\in K\setminus\{x,y\}$ With two technologies, the set of models becomes $M=\{(\mathcal{H}_{x},\mathcal{H}_{y}),(\mathcal{H}_{x},\mathcal{L}_{y}),(\mathcal{L}_{x},\mathcal{H}_{y}),(\mathcal{L}_{x},\mathcal{L}_{y})\}$, where subscripts are only used to keep track of the views corresponding to each technology. According to this notation, when Ann models technologies according to $(\mathcal{H}_{x},\mathcal{L}_{y})$, she holds the optimistic view about technology $x$ and the skeptical one for technology $y$.
The multi-technology setup admits many possible interpretations. For instance, our group could consist of board members of a company trying to decide which business project to prioritize between a number of promising, but uncertain alternatives. They could be investors deciding between two competing investment portfolios. They could be academics disagreeing over which research method, or theory, will yield the best answers to a research question. They might be collaborating artists, disagreeing over the best way to create an artwork or song. In all cases, Ann and Bob have access to alternative methods of production. Importantly, we maintain the assumption that the players’ interests are aligned, to some degree: each of them benefits if the other adopts the best production method — no matter how strong the disagreement over what such method is. [^28] For instance, if Ann and Bob are two scientists facing a similar research question, we assume that they share the common goal of pushing knowledge as far as possible in their field.
With two technologies, disagreement can be of two kinds. As in section 4, it can be “vertical,” when players differ in their degree of optimism. [^29] This is the case, for instance, if Ann holds $m^{A}=(\mathcal{H}_{x},\mathcal{H}_{y})$ while Bob holds $m^{B}=(\mathcal{L}_{x},\mathcal{L}_{y})$, or if $m^{A}=(\mathcal{H}_{x},\mathcal{H}_{y})$ and $m^{B}=(\mathcal{H}_{x},\mathcal{L}_{y})$. But it can also be “horizontal,” when they are optimistic about different technologies — so that they cannot be ranked in terms of their optimism. For instance, this happens if $m^{A}=(\mathcal{H}_{x},\mathcal{L}_{y})$ and $m^{B}=(\mathcal{L}_{x},\mathcal{H}_{y})$.
@@
The intuition for the result is as follows. If Ann and Bob start the game like-minded, in the best case scenario they will exert effort $e^{\mathcal{H}}$ in every period — this happens provided that they are optimistic about at least one technology. With initial horizontal disagreement, for any technology assignment, players are motivated to exert more effort in $t=1$ than statically optimal given the corresponding technology view. By working harder, they increase the probability of their colleague’s model change, which, they believe, would ensure that the latter adopts the best technology in $t=2$. Hence, when initially assigned to the technology they are optimistic about, both Ann and Bob choose effort levels above $e^{\mathcal{H}}$. When ($Q_{x}=Q_{y}$), the increase in effort always translates into higher expected output.
@@
    If the opposing models $(\mathcal{H}_{x},\mathcal{L}_{y})$ and $(\mathcal{L}_{x},\mathcal{H}_{y})$ are equally likely to be correct, a team with horizontal disagreement performs on average better than any like-minded team. Formally, if $Q=(\mathcal{H}_{x},\mathcal{L}_{y})$ with probability $p(\mathcal{H}_{x},\mathcal{L}_{y})=\frac{1}{2}$ and $Q=(\mathcal{L}_{x},\mathcal{H}_{y})$ with probability $p(\mathcal{L}_{x},\mathcal{H}_{y})=\frac{1}{2}$, then for each equilibrium profile $s\in\hat{S}$ it holds
    $\mathbb{E}_{p}\left[Y_{s}((\mathcal{H}_{x},\mathcal{L}_{y}),(\mathcal{L}_{x},\mathcal{H}_{y}),Q)\right]\geq\mathbb{E}_{p}\left[Y_{s}(m,m,Q)\right]\quad\forall m\in M.$
@@
A manager wants to form a team of two engineers to develop a new product. The product can be based on two very different innovative technologies. Some of the engineers of the firm are optimistic about both alternatives (i.e., hold view $\mathcal{H}$ for both technologies), others support only one. The manager knows that typically one approach proves better than the other — one technology is of type $\mathcal{H}$ and the other of type $\mathcal{L}$. — but, from her point of view, both technologies are equally promising ex-ante. Proposition 5.2 (ii) tells us that the manager will maximize the expected output of the team if she picks co-workers with opposite views, $(\mathcal{H},\mathcal{L})$ vs $(\mathcal{L},\mathcal{H})$, provided that $\mathcal{H}$ and $\mathcal{L}$ are equally falsifiable. The benefit of disagreement is twofold: first, both team members will try harder to make more progress during the first production period. Such progress is valuable on its own. Second, the higher first-period effort conveys additional information about the promise of the two technologies, making it more likely that the team will adopt the best technology in the second period.
The takeaway of this section is simple and yet, we believe, important and non-trivial. When two similarly good production technologies are available, disagreement on the best way to produce might increase group production. We focused on a mechanism that builds on the idea of different perspectives and shows that these differences can be useful even when they do not complement each other, that is, when there is competition of ideas. When they have the common goal of increasing each others’ output, disagreeing people will challenge each other, and work harder to prove, with their successes, that their perspective is valid, beneficial, and worth adopting. The resulting increase in group output is where — we believe — lies the “disagreement dividend.”
@@ 6 Informativeness, Externalities, and Welfare
A similarly crucial role is played by the assumption that each team member is ceteris paribus better off when other members produce more. The positive externality creates the incentive to persuade disagreeing co-workers to abandon skeptical views, as well as the one to convince them to adopt the best technology. Clearly, if other players’ output entered utility negatively, the incentive would be very different. To see this, consider equations 1, 2 and 3 of our illustrative example of section 2, but let $\beta<0$ — so that externalities are negative. By inspecting the equations, it is easy to see that $\beta<0$ implies that the value of triggering a co-worker’s change in mind at the end of the first period becomes negative: as a result, disagreement reduces Ann’s first-period effort and expected output, rather than increasing them.
@@
We omit the formal proofs for the claim, as such proofs would follow very closely the ones reported in the appendix. Instead, we provide the intuition for the one-technology case (the multiple-technology case follows the same intuition). Let the alternative assumptions of claim 1 hold, and consider a disagreeing team. Note that the player with view $\mathcal{H}$ will be hurt from a change of mind of the $\mathcal{L}$-view player: by assumption 2, the skeptic works harder after becoming optimist, damaging her co-worker through the negative production externalities of point (i). This force pushes the optimistic player to reduce information arrival in the first period, in order to decrease the (subjective) probability of a co-worker’s change in mind. However, when the relation between effort and informativeness is inverse — as of point (ii) — the way to reduce information arrival is by working harder. Hence, even under these alternative assumptions, disagreement pushes optimists to work harder and, if negative externalities are strong enough, a disagreeing team will produce on average more than any like-minded team. [^31] A few details of the propositions will indeed need to be intuitively modified for the result to hold under the assumptions of claim 1. In particular, the second part of proposition 4.2 (ii) holds for $\Delta$ negative enough; and the requirement for inequalities to be strict becomes that $\phi^{m^{-i}}_{m^{i}}(\mathbf{e},\mathbf{k}|m^{i})$ is strictly decreasing in effort. Regardless of this equivalence, we believe that our original specification — with positive externalities and more information arriving the more a team works on a project — is particularly realistic and captures well the zest of most of our motivating examples.
We conclude the discussion with a word on the welfare implications of disagreement. A welfare analysis would be particularly complex in our setup, which does not impose strong assumptions on the true process $Q$. Without such assumptions, it is hard to tell, for instance, whether the effort levels that maximize team members’ joint expected payoffs (or are Pareto efficient in terms of expected utility) are above or below $e^{\mathcal{H}}$ and $e^{\mathcal{L}}$. What we can say with certainty is that if the assumptions of proposition 5.2 part (i) hold and if, additionally, $Q_{x}=Q_{y}=\mathcal{H}$, any Pareto efficient stage effort must be above $e^{\mathcal{H}}$. Hence the boost in players’ effort generated by horizontal disagreement can — as in the example presented in section 2 — lead to a Pareto improvement. Not much can be concluded, however, in general. [^32] For instance, note that with $|K|=1$ and $Q=\mathcal{H}$, the expected payoff of an optimist is always higher when she’s paired with another optimist. In such case, disagreement cannot lead to a Pareto improvement in terms of expected payoffs.